OpenStudy (anonymous):

How do you multiply and divide integers? I asked the question before but lost the paper. Help?

OpenStudy (nowhereman):

Multiplication is done like: \[0\cdot n = 0,\quad (m+1)\cdot n = m\cdot n + n\]\[(-m)\cdot n = m\cdot (-n) = -(m\cdot n)\] and for division you have to find m: \[m = \frac k n ⇔ m\cdot n = k\]

OpenStudy (anonymous):

you mulitlpy th numbers then you lok at the sign same sign = positve different sign =negitve

OpenStudy (anonymous):

So a negativs and a negative equals a positive and a postive and a negative equals a negative???

OpenStudy (nowhereman):

yes